Hi, I'm Lauren and I really need help with this Math Problem. Please help me and e-mail me back as soon as possible! Thank you I really Apreciate it!
Circle the prime numbers.
2 7 60 71 82 93

Ps/ I dont even know what prime numbers are! Thanks Again!

Answer by checkley71(8403)

prime numbers are numbers that are not multiples of any other number ( except 1).
2 is prime because it s only factors are 1 & 2 (1*2=2)
7 is another prime number (1*7=7)
60 is not a prime number because it has many factors (2*30=60, 3*20=60, 4*15=60, 5*12=60 & 6*10=60)
71 is a prime number
82 is not prime (2*41=82)
93 is not a prime number (3*31=93)
NOTE: any number ending in an even number 2(exception is 2),4,6,8,0 are not prime. also numbers ending in 5 (except 5) are not prime.